18. Find the three 3rd roots of z= -4V2 + 4V2 i, show work. (18, wo =) (18, wi =) (18, W2 =)
Answer #1

ا ا ا ا ا ا ا ا م ن و ه ی 0.26 27 - 2017 - و 2 ۰۲ - ۱ / ۰۶۶= (5) اد + (5) جولته = 2 (ک -2 -2 (.+ 5.1747=-=J2 +۱۶۶ رود با تو